OpenStudy (anonymous):

solve the pair of simultaneous equations:y=8-x , 2x^2+xy=-16

OpenStudy (anonymous):

this only one equation? it's not a pair

OpenStudy (anonymous):

put y=8-x in the second equation: 2x^2+x(8-x)+16=0 then 2x^2-x^2+8x+16=0 x^2+8x+16=0 we see that x^2+8x+16=(x+4)^2 (formula (a+b)^2=a^2+2ab+b^2) so x is equal to -4 hence y is equal to 12